A rectangle has a base of 4x and a height of 3x + 5. If the perimeter is 150 units, what is the value of
Answer by ewatrrr(24785)

 
Hi,
.
2L + 2w = P
2(4x) + 2(3x+5) = 150
14x = 140
x = 10
base: 40units
Height: 35units
And...
2L + 2w = P
80units + 70units = 150units CHECKS